OpenStudy (anonymous):

Select the rule for the nth term of this arithmetic sequence. 4.3, 7, 9.7, 12.4, 15.1,...

OpenStudy (anonymous):

@Loser66 Can you help?

OpenStudy (loser66):

answer me 7-4.3 =?

OpenStudy (anonymous):

2.7

OpenStudy (loser66):

9.7-7=

OpenStudy (anonymous):

2.7

OpenStudy (loser66):

one more time, 12.4 -9.7=?

OpenStudy (anonymous):

Lol. 2.7

OpenStudy (loser66):

yeah that means the second term = the first term + 2.7, right? the third term = the second term + 2.7, right? the fourth term = the third term + 2.7 again, right? so on and so on. So the rule is \[A_n=A_{n-1}+2.7\]. do you want that rule or need another form? we can construct another form, but I don't know what you really want

OpenStudy (anonymous):

Oooooh. Well I need to find a term. Here are my choices: 4.3n + 2.7(n-1) 1.6 + 2.7n 4.3 + 2.7(n - 1) 4.3 + 2.7n

OpenStudy (loser66):

oh, I see. want to construct by yourself with my guide or just answer?

OpenStudy (anonymous):

Well I definitely want to learn how to do this, so I'd like to attempt it

OpenStudy (anonymous):

With a nudge in the right direction lol

OpenStudy (loser66):

ok, let n is the order of the term, so n =1 is the first term, right? n=1, A_1 = 4.3, got it?

OpenStudy (anonymous):

Ok. So 4.3 would really be the second term?

OpenStudy (loser66):